Consider an individual who earns $100,000, pays $7,000 in child care expenses for one child, pays the maximum ($26,010) in Registered Pension Plan (RPP) contributions, makes charitable donations of $200 (15 percent credit), and pays $20,000 in medical expenses. Medical expenses in excess of 3 percent of net income are deductible at 20 percent. The statutory marginal tax rate is 17 percent. What is the federal tax payable? What is the individual’s actual (effective) marginal tax rate? [6 marks]
1
Expert's answer
2019-06-11T10:24:03-0400

Taxable income: $100,000-$7,000-$26,010-$20,000*0.8=$50,990

Gross tax:$50,990*0.17=$8,668

Tax credits:$200*0.15=$30

Tax payable:$8,668-$30=$8,638

Effective tax rate:$8,638/$50,990=16.94%